Sam is three times as old as Steve. In 8 years, he will be twice as old.
Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52792)

.
Thx for letting us know.


Answer by Edwin McCravy(20056)

Obviously you wanted to know the age of each.  Ikleyn knew that was what
you wanted.  She just wanted to impress upon you to be careful to
state the whole problem, not just part of it.

Let X = Sam's age
Let Y = Steve's age.

Sam is three times as old as Steve.
X = 3Y

In 8 years, he (Sam) will be twice as old (as Steve).
Sam's age in 8 years = X+8
Steve's age in 8 years = Y+8

X+8 = 3(Y+8)

The system of equations is

X = 3Y
X+8 = 2(Y+8)

Substitute 3Y for X in:   X+8 = 2(Y+8)

3Y+8 = 2(Y+8)

Solve for Y. That will be Steve's age.
Then multiply that by 3 to get Sam's age.

Edwin